mr c is twice as old as his son,in 42 years the ratio of their ages will be4:3 what is the sons age

Answer by checkley77(12844)

ages are x for the son & 2x for the mr. now
in 42 years they will be x+42 for the son & 2x+42 for the mr.
4/3=(2x+42)/(x+42)
3(2x+42)=4(x+42)
6x+126=4x+168
6x-4x=168-126
2x=42
x=42/2
x=21 is the age of the son now.
Proof:
4/3=(2*21+42)/21+42)
4/3=(42+42)/63
4/3=84/63
3*84=4*63
252=252